OpenStudy (anonymous):

FAN AND MEDAL: NEED HELP PLEASE! Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (anonymous):

@PinkSapphire @Jagking

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well if sally received 25% of the vote Maya received 75% of the vote so if the total number of votes are 125 000 then Sally = 125000 x 25/100 Maya=125000 x 75/100 hope it helps

OpenStudy (anonymous):

It helps a lot...Do you mind helping me with some other word problems like this one? @campbell_st

OpenStudy (anonymous):

Or can somebody else help me? @PinkSapphire @Jagking

OpenStudy (anonymous):

WHTS UP?

OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (anonymous):

We are working on Two-Variable Equations...

OpenStudy (pinksapphire):

R=250

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

x=250

OpenStudy (anonymous):

dang. u always bet me pinky.

OpenStudy (pinksapphire):

LOL

OpenStudy (pinksapphire):

I will explain...

OpenStudy (anonymous):

okay. thank you.

OpenStudy (anonymous):

u take 1000 and mutiply it wit 700. which will give u 700000. then u subtract tht wit 1,000,000. then u divide tht by 1200. then thts how u get r.

OpenStudy (pinksapphire):

700A+1200R=1000000 700(1000)+1200R=1000000 700000+1200R=1000000 -700000 -700000 1200R=300000 1200R/1200=300000/1200 R=250

OpenStudy (anonymous):

ha. beat u this time.

OpenStudy (anonymous):

Thank you both... A LOT! I have so many to do and they are so confusing!

OpenStudy (pinksapphire):

Do you need any more help?

OpenStudy (anonymous):

If you don't mind. Can I open it in another question though?

OpenStudy (pinksapphire):

Sure.

OpenStudy (anonymous):

yea. bring it. i'm beaten u this time pinky.

OpenStudy (pinksapphire):

Yeah you did:(